### v1

# CasualHDRSplat: Robust High Dynamic Range 3D Gaussian Splatting from Casually Captured Videos

Abstract
Recently, photo-realistic novel view synthesis from multi-view images, such as neural radiance field (NeRF) and 3D Gaussian Splatting (3DGS), have garnered widespread attention due to their superior performance. However, most works rely on low dynamic range (LDR) images, which limits the capturing of richer scene details. Some prior works have focused on high dynamic range (HDR) scene reconstruction, typically require capturing of multi-view sharp images with different exposure times at fixed camera positions during exposure times, which is time-consuming and challenging in practice. For a more flexible data acquisition, we propose a one-stage method: CasualHDR to easily and robustly reconstruct the 3D HDR scene from casually captured videos with auto-exposure enabled, even in the presence of severe motion blur and varying unknown exposure time. CasualHDR contains a unified differentiable physical imaging model which first applies continuous-time trajectory constraint to imaging process so that we can jointly optimize exposure time, camera response function (CRF), camera poses, and sharp 3D HDR scene. Extensive experiments demonstrate that our approach outperforms existing methods in terms of robustness and rendering quality.
https://github.com/WU-CVGL/CasualHDRSplat
Figure 1: a) Our method can reconstruct 3D HDR scenes from videos casually captured with auto-exposure enabled. b) Our approach achieves superior rendering quality compared to methods like Gaussian-W and HDR-Plenoxels. c) After 3D HDR reconstruction, we can not only synthesize novel view, but also perform various downstream tasks, such as 1) HDR exposure editing, 2) Image deblurring.
11footnotetext: Equal contribution.
1 Introduction
Photo-realistic Novel View Synthesis (NVS) are vital areas in computer vision with applications in VR/AR, autonomous driving, and embodied AI, offering immersive experiences for both humans and AI agents. Neural Radiance Fields (NeRFs)[36] and 3D Gaussian Splatting (3DGS)[18], through reconstructing radiance fields, have become mainstream methods in the NVS field due to their high-quality rendering, inspiring a large number of subsequent research works[40, 64, 54, 13].
However, most NVS methods struggle with high-contrast inputs, assuming good-quality images with consistent exposure conditions and low dynamic range (LDR). On the other hand, limited dynamic range of the inputs hinders 3D scene representations from reconstructing fine details in high dynamic range (HDR) environments, thereby restricting its further applications. Although 2D HDR contents have been standardized, consumed and exploited in recent years [12, 15, 1], 3D HDR free-viewpoint (volumetric) content is still a new concept with great potential. Therefore, reconstructing 3D high dynamic range (HDR) scenes is of significant practical value for achieving better visual effects and meeting the needs of downstream tasks.
Current 3D HDR NVS methods can be divided into two categories. The first category, e.g. RawNeRF[37] and LE3D[16] etc., takes in noisy RAW images, aiming to reconstruct noise-free 3D HDR scenes. The second category, represented by [13, 4, 17, 53, 58], draws inspiration from HDR imaging (HDRI), using multi-exposure LDR images as inputs to reconstruct the 3D HDR scene while learning camera response function (CRF). However, the strict inputs limit their flexibility and broader applications. The challenges include: 1) Data acquisition of RAW images and accurate exposure time is usually expensive due to the use of professional equipment; 2) In low-light conditions, long exposure times increase the risk of motion blur from camera shake, reducing reconstruction quality; 3) The geometric consistency will be compromised if given inaccurate camera pose initialization, as the camera poses are not being optimized. Thus, a key challenge is reducing the cost of data acquisition, enabling high-quality and robust 3D HDR scene reconstruction with consumer-grade devices.
Most modern consumer-grade cameras use auto-exposure during video recording, automatically adjusting exposure time based on ambient lighting. This expands the captured dynamic range in the video, making it possible to reconstruct 3D HDR scenes. However, naively applying these videos to existing HDR 3D reconstruction methods presents several challenges: 1) Although some cameras support providing exposure time in images stored in EXIF metadata, it is often impossible to obtain the exposure time for each frame in a video format; 2) Auto-exposure can cause inconsistencies in brightness between frames, leading to pose estimation errors in structure from motion (SfM) frameworks; 3) Videos suffer from severe motion blur due to the camera movement during exposure time.
To address these challenges, we cannot assume that the camera is static during the exposure time, as done in previous methods. Instead, we must also account for camera motion during this period. By analyzing the physical imaging process, we found that motion blur and image brightness variations are both directly related to the exposure time: a longer exposure time typically leads to more severe motion blur and higher image brightness. In previous works [13, 54], it has been shown that camera motion blur and image brightness can be decoupled as camera movement and the accumulation of irradiance during the exposure time. Thus, camera motion blur and varying image brightness can be constrained by the exposure time.
Building on above reasoning, we first introduced the continuous-time trajectory to represent the camera motion during video recording rather than addressing each camera pose as previous works[54, 64], so that the camera poses and exposure times can be jointly optimized. Subsequently, we design a camera response function (CRF) module which can convert the accumulation of irradiance (HDR image) over the exposure time to image brightness (LDR image). As mentioned above, exposure time links motion blur and image brightness. We incorporate camera motion into the physical imaging model and build a unified differentiable rendering framework, allowing the trajectory, CRF, and exposure time to be jointly optimized and constrained.
With explorations above, we propose a one-stage method called CasualHDRSplat, a unified 3D HDR reconstruction framework that couples the physical imaging model with camera motion, improving the robustness and flexibility. In our designed unified imaging model, the continuous-time camera trajectory on $\mathbb{SE}(3)$, exposure time, and camera response function are jointly optimized and mutually constrained. Therefore, our approach doesn’t require ground truth exposure times as inputs like previous methods.
In summary, our contributions are as follows:
  • CasualHDRSplat, a unified imaging model that first applies continuous-time trajectory to constrain physical imaging model so that camera poses, CRF, exposure times and 3D HDR representation can be jointly optimizes, enabling users to reconstruct 3D HDR scenes from casually captured videos at a low cost.
  • A dataset that includes both synthetic data and real-world data, where each video contains severe variations in brightness and camera motion blur, that can be useful to the community to further investigate into this problem.
  • With extensive experiments, we demonstrate how to utilize this model to reconstruct high-quality HDR scenes from casual videos, and exhibit state-of-the-art performance across all datasets.
2 Related Work
2.1 High Dynamic Range Imaging
High Dynamic Range Imaging (HDRI) enhances luminosity beyond standard digital imaging by merging multi-exposure LDR images. In video capture, alternating long and short exposures achieves similar effects. Recently, deep learning approaches have treated HDRI as an image domian translation task, designing networks to convert LDR to HDR images. However, camera motion during exposure time often lead to ghosting artifacts. To address this, [11] proposed an adaptive metering algorithm to adjust exposure and reduce motion artifacts, while other methods use spatial attention to mitigate motion blur. With the advent of 3D scene representations such as NeRF and 3DGS, methods like [13, 17, 4, 14, 53] have emerged to reconstruct 3D HDR scenes and calibrate CRFs simultaneously. While these methods are effective, they often rely on precise exposure times and struggle with motion blur, highlighting the need for improved robustness.
2.2 Image deblurring
Image deblurring aims to restore sharp images from blurred ones and can be categorized into three types. The first type uses hand-crafted priors, like total variation and heavy-tailed gradient priors, to solve for the blur kernel [21, 6], but struggles with different kernels producing similar effects. The second type, deep learning-based, achieves end-to-end restoration using large datasets [61, 51]. The third type leverages multi-view blurred images to reconstruct the 3D scene and deblur the images with geometric constraints. Works like [34, 42, 22] jointly learn the blur kernel and radiance field, while BAD-NeRF[54] introduces a physical motion blur model that jointly optimizes the radiance field and camera trajectories. Following BAD-NeRF[54], emerging works [23, 26, 25, 24, 49, 5, 41, 64, 60, 27, 43, 50] have demonstrated the effectiveness of using $\mathbb{SE}(3)$ trajectory representations for joint 3D reconstruction and blur image recovery.
2.3 Robust Novel View Synthesis
Novel view synthesis aims to generate images from arbitrary viewpoints using input images with known poses. Neural Radiance Fields (NeRF)[35] and 3D Gaussian Splatting (3DGS)[18] advanced this field by reconstructing 3D scene as radiance field and achieve great success.
Most synthesis methods assume high-quality input data, but quality degrades significantly with blurry images, large exposure variations, or inaccurate poses. NeRF-W[35] and Gaussian-W[62] address these challenges by adding optimizable appearance vectors to model varied conditions. HDR-NeRF[13] and HDR-Plenoxels[17] reconstruct HDR scenes even with dynamic exposure variations. Other methods, like [8], handle pose inaccuracies, while BAD-Gaussians[64] incorporates camera motion models to deblur inputs during reconstruction.$I^{2}$-SLAM[2] is a concurrent approach capable of processing images with exposure inconsistencies and blur, though it focuses on RGB-D SLAM and has different representations of trajectory and CRF. These methods collectively improve robustness for NVS under challenging conditions.
3 Method
In this section, we will provide a detailed explanation of our proposed CasualHDRSplat, which takes video captured with auto-exposure enabled as input. We first use continuous-time trajectory to represent the camera motion during video recording rather than addressing each camera pose individually, so that the camera pose and exposure time can be jointly optimized through bundle adjustment. Subsequently, we design a camera response function (CRF) module which can convert the accumulation of irradiance (HDR image) over the exposure time to pixel value (LDR image). Due to the fact that the exposure time is the connection between the continuous-time trajectory and CRF module, we can treat them as a unified differentiable imaging model thus the trajectory, CRF and exposure time can be jointly optimized and constrained together.
3.1 Preliminary: 3D Gaussian Splatting
3DGS represents the scene as 3D Gaussian primitives denoted as $\mathbf{G}$. Each primitive is characterized by a mean position $\bm{\mu}\in\mathbb{R}^{3}$, opacity $\mathbf{o}\in\mathbb{R}$, color $\mathbf{c}\in\mathbb{R}^{3}$, and a 3D covariance matrix $\mathbf{\Sigma}\in\mathbb{R}^{3\times 3}$. To ensure $\mathbf{\Sigma}$ remains positive semi-definite, it is parameterized by a scaling matrix $\mathbf{S}\in\mathbb{R}^{3}$ and a rotation matrix $\mathbf{R}\in\mathbb{R}^{3\times 3}$ stored as a quaternion $\mathbf{q}\in\mathbb{R}^{4}$. During rendering, the 3D Gaussians are projected onto the image plane at a given pose $\mathbf{P}_{i}=\{\mathbf{R}_{c}\in\mathbb{R}^{3\times 3},\mathbf{t}_{c}\in\mathbb{R}^{3}\}$ , transforming $\mathbf{\Sigma}$ into a 2D covariance matrix $\mathbf{\Sigma^{\prime}}\in\mathbb{R}^{2\times 2}$. These can be mathematically expressed as:
$\mathbf{G}(\mathbf{x})=e^{-\frac{1}{2}(\mathbf{x}-\bm{\mu})^{\top}\mathbf{\Sigma}^{-1}(\mathbf{x}-\bm{\mu})}$ (1)
$\mathbf{\Sigma}=\mathbf{R}\mathbf{S}\mathbf{S}^{T}\mathbf{R}^{T},\mathbf{\Sigma^{\prime}}=\mathbf{J}\mathbf{R}_{c}\mathbf{\Sigma}\mathbf{R}_{c}^{T}\mathbf{J}^{T},$ (2)
where $\mathbf{J}\in\mathbb{R}^{2\times 3}$ is the Jacobian of the affine approximation of the projective transformation. Next, the 2D Gaussians undergo depth sorting followed by tile-based rasterization. The final color values for individual pixels are obtained using $\alpha$-blending:
$\mathbf{C}(x,y,\mathbf{P}_{i})=\sum_{i=1}^{N}\mathbf{c}_{i}\alpha_{i}\prod_{j=1}^{i-1}(1-\alpha_{j})$ (3)
$\alpha_{i}=\mathbf{o}_{i}\cdot\exp(-\sigma_{i}),\quad\sigma_{i}=\frac{1}{2}\Delta_{i}^{T}\mathbf{\Sigma^{\prime}}^{-1}\Delta_{i},$ (4)
where $\mathbf{c}_{i}$ is the learnable color of each Gaussian, and $\alpha_{i}$ is the alpha value determined by the 2D covariance $\mathbf{\Sigma^{\prime}}$ multiplied by the learned opacity $\mathbf{o}$. ${\rm\Delta}_{i}\in\mathbb{R}^{2}$ represents the offset between the pixel center and the 2D Gaussian center. The above derivations show that the rendered pixel color, $\mathbf{C}$ in Eq.(3), is differentiable with respect to all learnable parameters $\mathbf{G}$ and camera poses $\mathbf{P}_{i}$, which is crucial for our bundle adjustment formulation.
Figure 2: The pipeline of CasualHDRSplat. Given a casually captured video with auto exposure, camera motion blur, and significant exposure time changes, we train 3DGS to reconstruct an HDR scene. We design a unified model based on the physical image formation process, integrating camera motion blur and exposure-induced brightness variations. This allows for the joint estimation of camera motion, exposure time, and camera response curve while reconstructing the HDR scene. After training, our method can sharpen the train images and render HDR and LDR images from specified poses.
3.2 Continuous Trajectory Representation
Cumulative $\mathbb{SE}(3)$ B-spline is a widely used continuous-time trajectory representation in robotics, especially in state estimation, sensor fusion and path planning [9, 33, 3, 45, 39, 10]. $\mathbb{SE}(3)$ B-spline has many excellent characteristics: such as $C^{2}$ continuity, locality and convex hull property that delicately incorporates gradient information and dynamic constraints, which converges quickly to generate smooth and feasible trajectories[65]. $\mathbb{SE}(3)$ B-spline allows for the calculation of pose, velocity, and accelerations at any timestamp given along a trajectory.
Targeting at unordered inputs, existing multi-view deblurring methods following [54] model the camera motion and estimate short splines for each frame separately, thus cannot utilize the cross-frame motion constrains and priors, given a continuous video as their input. Some methods [52, 28, 49, 29, 30, 47, 55] utilize basis functions to regularize continuous-time deformations to reconstruct dynamic scenes, but robust reconstruction from casual videos with a continuous-time camera trajectory representation has not yet been explored. To this end, this paper estimates the camera motion across the whole video with a continuous-time cumulative $\mathbb{SE}(3)$ B-spline trajectory.
Following [33], given a series of temporally uniformly distributed control knots, the pose $\mathbf{P}(t)$ at a given timestamp $t$ can be interpolated with 4 adjacent control knots, denoted as $\mathbf{T}_{\mathrm{0}}$, $\mathbf{T}_{1}$, $\mathbf{T}_{2}$ and $\mathbf{T}_{3}\in\mathbb{SE}(3)$:
$\mathbf{P}(t)=\mathbf{T}_{0}\cdot\prod_{j=0}^{2}\mathrm{exp}(\tilde{\mathbf{B}}(u)_{j+1}\cdot\bm{\Omega}_{j}),$ (5)
$\tilde{\mathbf{B}}(u)=\mathbf{C}\begin{bmatrix}1\\ u\\ u^{2}\\ u^{3}\end{bmatrix},\quad\mathbf{C}=\frac{1}{6}\begin{bmatrix}6&0&0&0\\ 5&3&-3&1\\ 1&3&3&-2\\ 0&0&0&1\end{bmatrix}.$ (6)
where $\tau$ represents the spline sampling interval, $u=\frac{t}{\tau}$, and $u$ lies within the interval $[0,1)$; $\tilde{\mathbf{B}}(u)_{j+1}$ denotes the $(j+1)^{th}$ element of the vector $\tilde{\mathbf{B}}(u)$, $\bm{\Omega}_{j}=\mathrm{log}(\mathbf{T}_{j}^{-1}\cdot\mathbf{T}_{j+1})$, based on the [44].
3.3 Physical Image Formation Model
The physical image formation process refers to a digital camera collecting scene irradiance during the exposure time $\Delta t$ and converting them into measurable electric charges, which are ultimately mapped into pixel values through the camera response function (CRF) defined by $F$. Assuming the camera moves along a continuous trajectory $t$ $\mapsto$ $\mathbf{P}(t)$ during exposure time $\Delta t$ with constant velocity, this process can be mathematically modeled as follows:
$\mathbf{B}(x,y)=F\left(\int_{t_{b}}^{t_{b}+\Delta t}\mathbf{H}\left(x,y,\mathbf{P}(t)\right)\mathrm{dt}\right)$ (7)
where $\mathbf{B}(x,y)\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ denotes the real captured image, $x,y\in\mathbb{R}^{2}$ represents the pixel location, $t_{b}$ denotes the timestamp when the shutter opens, $\mathbf{H}\left(x,y,\mathbf{P}(t)\right)$ represents scene irradiance mapped into camera at pose $\mathbf{P}(t)$ which is interpolated from the continuous trajectory. Additionally, if the camera moves during the exposure time, the camera will collect irradiance from different scene points, resulting in camera motion blur. The integral part in Eq. (7) can be discretized as follows:
$\mathbf{H}(x,y)\approx\frac{1}{N}\sum_{k=0}^{N-1}\mathbf{H}_{k}\left(x,y,\mathbf{P}(t_{k})\right)\Delta t$ (8)
$\mathbf{H}(x,y)\in\mathbb{R}^{\mathrm{H}\times\mathrm{W}\times 3}$ denotes blur HDR image, N represents the number of virtual latent sharp images, $\Delta t_{k}$ represents the exposure time of virtual camera $k$ and can be set as a constant equal to $\frac{\Delta t}{N}$, $t_{k}$ denotes the timestamp corresponding to virtual camera $k$, it can be calculated as $t_{b}+\frac{\Delta t}{N}*k$.
After obtaining $\mathbf{H}(x,y)$, we need to use the camera response function $F$, which includes image-varying white balance $\mathrm{WB}$ and tone mapping $\mathrm{TM}$, to convert it into an LDR image:
$\mathbf{B}(x,y)=\mathbf{F}(\mathbf{H}(x,y))=\text{TM}\circ\text{WB}(\mathbf{H}(x,y))$ (9)
$\text{WB}(\mathbf{c})=\begin{bmatrix}wb_{r},wb_{g},wb_{b}\end{bmatrix}^{T}\odot\begin{bmatrix}c_{r},c_{g},c_{b}\end{bmatrix}^{T}$ (10)
Due to the fact that RGB channels have different camera response curves for $\mathrm{TM}$, we adopt separate MLP for each channel. Unlike prior methods, we treat $\Delta t$ as an optimizable quantity rather than a precisely known parameter. Initially, $\Delta t$ can be assigned a random value. Since the exposure time directly affects the brightness and motion blur of the image, it will be gradually optimized to the actual value during the subsequent deblurring and HDRI processes. This significantly reduces the dependency on the exposure time and enhances the robustness of 3D HDR reconstruction.
3.4 Loss Function
Given a series of video frames moving along a continuous trajectory, we can estimate the learnable Gaussian primitives, the camera trajectory parameters, implicit CRF representation and the exposure time for each image. This estimation can be achieved by minimizing a loss function, which can be specifically expressed as follows:
$\mathcal{L}=\mathcal{L}_{\text{rec}}+\lambda_{\text{exp}}\mathcal{L}_{\text{exp}},\quad\mathcal{L}_{\text{rec}}=(1-\lambda)\mathcal{L}_{1}+\lambda\mathcal{L}_{\text{D-SSIM}},$ (11)
where $\mathcal{L}_{\text{rec}}$ can constrain the consistency between the rendered image $\mathbf{C}_{k}(\mathbf{x})$ (the $k^{th}$ blurry LDR image synthesized from 3D-GS using the aforementioned image formation model (Eq. 8)) and the input LDR image $\mathbf{C}^{gt}_{k}(\mathbf{x})$.
To accurately model significant exposure variations in the input images, the second term of the loss function normalizes the images to a medium exposure by scaling pixel intensities before computing discrepancies [32, 53]:
$\small\mathcal{L}_{\text{exp}}=\mathcal{L}_{1}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right)+\mathcal{L}_{\text{D-SSIM}}\left(\frac{\mathbf{C}^{gt}_{k}(\mathbf{x})}{\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})},\frac{\mathbf{C}_{k}(\mathbf{x})}{\bar{\mathbf{C}}_{k}(\mathbf{x})}\right),$ (12)
where $\bar{\mathbf{C}}^{gt}_{k}(\mathbf{x})$ and $\bar{\mathbf{C}}_{k}(\mathbf{x})$ represent the average pixel value of $\mathbf{C}^{gt}_{k}(\mathbf{x})$ and $\mathbf{C}_{k}(\mathbf{x})$. We set $\lambda_{\text{exp}}=0.25$ in all our experiments[20].
4 Experiments
BAD-Gaussians HDR-Plenoxels Gaussian-W Ours-random Ours-gt Reference
Figure 3: Qualitative comparison on the Girls-vicon sequence of the Realsense dataset in terms of NVS.
BAD-Gaussians HDR-Plenoxels Gaussian-W Ours-random Ours-gt Reference
Figure 4: Qualitative comparison on the Building sequence of the Smartphone dataset in terms of NVS.
BAD-Gaussians HDR-Plenoxel Gaussian-W Ours-random Ours-gt Reference
Figure 5: Qualitative comparison on the Trolley sequence of the synthetic dataset in terms of NVS.
Figure 6: Qualitative comparison on the Pool sequence of the synthetic dataset under training view.
4.1 Datasets
Synthetic datasets. We generated a synthetic dataset using Blender 3.6 with the Cycles engine, featuring four distinct scenes: Factory, Pool, Cozyroom, and Trolley. Each scene contains 77 images. The dataset generation combines physical motion blur imaging model[54] and tone mapping[13] from HDR to LDR. For each scene, images were assigned random exposure times, and captured using a continuous camera trajectory, generating sharp HDR images, which were averaged over the exposure period to create motion-blurred images.
Real datasets. Current related datasets use multi-view images with known exposure times from fixed viewpoints during exposure time, unlike our approach using casual videos for HDR reconstruction. Therefore, we collected a challenging real-world dataset, CasualVideo, using Intel RealSense D455 and Google Pixel 8 Pro on a DJI RS3 Mini gimbal. The dataset includes two subsets: RealSense and Smartphone. RealSense has four sequences (Yakitori, Toufu, Toufu-vicon, Girls-vicon), with ground truth poses for the latter two from Vicon. Smartphone has two sequences: Building and Fish. Since RealSense cannot provide current exposure times when auto-exposure is enabled, which limits baseline methods relying on exposure data, we implemented custom auto-exposure control with fixed aperture and ISO, following [48], on both cameras. To further evaluate our method on real-world data, we used the publicly available ScanNet[7], which features scenes recorded with auto-exposure enabled[2].
Table 1: Quantitative comparisons on the synthetic datasets in terms of novel view
Factory Pool Trolley Cozyroom
[…]
gsplat [18] 15.14 0.75 0.25 11.73 0.65 0.30 14.48 0.62 0.32 13.86 0.76 0.22
Gaussian-W [62] 23.68 0.75 0.26 23.28 0.69 0.62 17.83 0.64 0.34 27.16 0.85 0.15
BAD-Gaussians [64] 14.99 0.81 0.25 25.09 0.72 0.30 16.12 0.66 0.22 17.12 0.75 0.20
HDR-Plenoxels [17] 24.36 0.72 0.29 30.84 0.81 0.33 17.05 0.55 0.42 28.13 0.81 0.13
HDR-NeRF [13] 14.57 0.31 0.68 - - - - - - 13.62 0.32 0.77
CasualHDRSplat-random (ours) 30.25 0.89 0.10 32.63 0.91 0.09 25.14 0.81 0.24 29.62 0.86 0.10
CasualHDRSplat-gt (ours) 30.75 0.90 0.09 32.36 0.92 0.08 25.85 0.88 0.11 31.32 0.92 0.09
Table 2: Quantitative comparisons on the real-world datasets in terms of novel view.
Fish-pixel8pro Building-pixel8pro Toufu-vicon Girls-vicon
[…]
gsplat 23.20 0.82 0.16 25.99 0.81 0.11 24.34 0.81 0.28 23.81 0.77 0.28
BAD-Gaussians 24.28 0.78 0.14 26.93 0.82 0.11 24.22 0.82 0.24 23.95 0.77 0.28
HDR-Plenoxels 19.39 0.53 0.65 26.87 0.81 0.15 17.90 0.51 0.69 26.73 0.84 0.30
Gaussian-W 26.13 0.83 0.15 27.99 0.82 0.11 26.38 0.83 0.29 26.88 0.86 0.25
CasualHDRSplat-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
CasualHDRSplat-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
4.2 Implementation Details
We implemented our method in PyTorch using the gsplat[59] with MCMC strategy[19]. The optimization of all the parameters was done using the Adam optimizer, keeping the Gaussian primitive learning rate consistent with gsplat. To balance performance and efficiency, we set the number of virtual camera poses (n in Eq. 8) to 10.
For initialization, we used HLoc[46] on our Sythetic and Realsense dataset .For the ScanNet[7] and our Smartphone datasets, we used DPV-SLAM [31] since HLoc failed due to poor image quality. We use NVIDIA RTX 3090 on synthetic data and A6000 on real data, respectively.
4.3 Baseline Methods and evaluation metrics
To evaluate the robustness of our method in learning accurate scenes representation under poorly exposed conditions and server motion blur, we compared it against scene reconstruction methods that handle brightness variations, e.g. HDR-NeRF[13], HDR-Plenoxels[17], Gaussian-W[62], as well as method for scene reconstruction from blurred images, such as BAD-Gaussians[64]. In addition, 3D-GS[18] implemented by gsplat[59] was included as the comparison baseline. Following previous works[17, 4, 13], we evaluate LDR images rendered from the learned scene used metrics such as PSNR, SSIM[57], and LPIPS[63]. Furthermore, to evaluate whether our method effectively recovers camera motion trajectories, we compared it with pose estimation method, e.g. HLoc[46], DPV-SLAM[31], BAD-Gaussians[64]. For pose estimation accuracy, we utilize absolute trajectory error (ATE) with mean and std as the metric. Since $I^{2}$-SLAM[2] and HDR-GS[4] are not open-sourced, we cannot compare our method against them.
4.4 Quantitative evaluation results.
We conducted experiments with two settings: randomly initialized exposure times (CasualHDRSplat-random) and ground truth exposure times (CasualHDRSplat-gt). We evaluated our method’s performance on novel view synthesis, image deblurring, and pose estimation (ATE) tasks. More results on the ScanNet dataset and two Realsense scenes can be found in the supplementary materials.
Due to the blurry nature of most real-world images, we used 5 to 10 sharp images per sequence for evaluation. As shown in Table 1 and Table 2, our method significantly outperforms previous methods in novel view synthesis. CasualHDRSplat-random, even with randomly initialized exposure times, surpasses prior works by jointly optimizing exposure times and CRF representation. Unlike HDR-NeRF[13], our method reconstructs HDR scenes accurately from degraded images without measured exposure times, whereas HDR-NeRF fails on all real dataset scenes.
By modeling the physical principles of camera imaging, our method also outperforms HDR-Plenoxels[17] and Gaussian-W[62]. Moreover, our approach leverages spline representations for optimizing camera motion trajectories, enabling effective scene learning, which other methods struggle with in the absence of ground truth poses.
Table 3: Ablation studies on the $ratio$ for initializing camera motion spline.
Pool Factory
[…]
0.5 29.89 0.83 0.12 23.25 0.69 0.20
1.0 30.49 0.83 0.11 23.93 0.70 0.16
1.5 31.13 0.84 0.10 24.78 0.74 0.16
2.0 32.01 0.87 0.10 25.64 0.77 0.16
2.5 32.04 0.88 0.10 26.90 0.81 0.14
3.0 32.95 0.88 0.10 27.25 0.84 0.15
3.5 33.13 0.89 0.09 27.54 0.84 0.14
4.0 33.63 0.90 0.08 27.60 0.84 0.14
Table 4: Ablation study: Average results across two datasets to investigate the effect on model performance.
Deblur Exp. Opt. CRF Conti. Traj. Average
[…]
14.50 0.755 0.235
16.06 0.78 0.225
20.04 0.665 0.29
19.98 0.675 0.29
26.03 0.79 0.12
28.43 0.84 0.10
Table 5: Quantitative comparisons on the synthetic datasets in terms of deblur
Factory Pool Trolley Cozyroom
[…]
BAD-Gaussians[64] 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
CasualHDRSplat-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
CasualHDRSplat-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
Table 6: Quantitative comparisons for pose estimation on the Realsense sequences with Vicon motion captured groundtruth. The results are in the absolute trajectory error metric (ATE) with units in centimeters.
HLoc DPV-SLAM BAD-GS CasualHDRSplat-random CasualHDRSplat-gt
Toufu-vicon .4644$\pm$.3921 .4043$\pm$.3877 .3935$\pm$.4212 .3687$\pm$.3874 .3595$\pm$.3462
Girls-vicon $1$9557$\pm$.8231 .8548$\pm$.8628 .8294$\pm$.8834 .6478$\pm$.8268
Table 5 shows that our method achieves superior performance in image deblurring task compared to BAD-Gaussians[64]. This is because our method can recover accurate scene representation from images affected by both motion blur and poor exposure.
The experimental results presented in Table 6 demonstrate that our method outperforms prior approaches in the pose estimation task. HLoc, which relies on feature point matching, exhibits poor performance under conditions of varying brightness and motion blur. Although DPV-SLAM[31] and BAD-Gaussians[64] can operate effectively in the presence of motion blur, they struggle to tolerate environments with high-contrast and varying exposure time. This indicates that our method can robustly estimate continuous camera trajectories under high-contrast environments, within varying exposure time and motion blur.
4.5 Qualitative evaluation results.
The qualitative comparisons of the NVS and deblurring tasks on both synthetic and real datasets are shown in Figure 3, 4, 5 and 6. The experimental results indicate that our method outperforms previous approaches. This demonstrates that our method can effectively learn scene representations from images that simultaneously exhibit varying exposure time and motion blur, while prior works lack robustness given the challenging conditions and failed to reconstruct high-quality 3D HDR scene. Note that BAD-Gaussians can deblur the training views as ours, however, due to its failure of pose optimization, the rendered images are misaligned with others.
4.6 Ablation studies.
Initialization for camera motion spline. In our method, the camera motion spline needs to be initialized by leveraging the poses estimated from HLoc [46] or DPV-SLAM [31]. Therefore, the configuration of initialization will impact the performance of our method. We define a ratio representing the number of control knots of spline divided by the number of input images and evaluate the effect of the ratio. The results in Table 4 indicate that model performance improves until it saturates as the ratio increases. We set $\textit{ratio}=3.0$ for all experiments to ensure a trade-off between the performance and computational overhead.
Effect of each module. Deblur represents the method’s ability to remove blur, Exp. Opt. indicates exposure time optimization, CRF represents whether the model includes a CRF module, and Conti. Traj. refers to the use of continuous trajectories to represent camera motion. The results presented in Table 4 highlight several key findings: 1) Utilizing splines to represent the continuous camera trajectory significantly enhances model performance, achieving approximately a 24% improvement in PSNR. 2) Jointly optimizing exposure time while learning an implicit representation of the CRF substantially boosts performance, leading to a 42% increase in PSNR. This demonstrates that our method can robustly reconstruct HDR scenes in environments with varying brightness. 3) Representing motion blur as the average of a series of sharp images over the exposure time yields a 9% improvement in PSNR, showing that our approach effectively handles input images with motion blur. In summary, the proposed representation of continuous trajectories and the joint optimization of exposure time with CRF contribute significantly to the model’s performance.
5 Conclusion
In this paper, we introduce a novel method CasualHDRSplat for reconstructing 3D HDR scenes from casually captured videos, which often exhibit limited dynamic range and motion blur. Our method reconstruct 3D HDR scene and generate LDR images given specified exposures and camera poses, providing high robustness and flexibility. By leveraging the camera auto-exposure, we incorporate the latent high dynamic range of captured videos into a unified physical imaging model. This allows for the joint optimization of exposure time, continous-time camera trajectory, and CRF, enabling accurate 3D HDR scene reconstruction. Extensive experiments demonstrate that our method outperforms previous approaches in 3D HDR reconstruction.
[…]
Supplementary Material
6 Appendix
In the supplementary material, we present more quantitative and qualitative experimental results for image rendering under both training and novel viewpoints. We also visualized the results of camera motion estimation and performed a qualitative comparison. The rendered novel view high frame-rate HDR video is presented in the supplementary video. We will present each part as follows.
6.1 HDR editing
Figure 7: Qualitative results of HDR editing with various exposure times.
The results in Figure 7 demonstrate that our method can accurately learn HDR scenes and the brightness of the rendered images can be adjusted by manually changing the exposure time.
6.2 More experimental results of exposure time estimation.
Table 7: Quantitative results for exposure time estimation.
Scene Pearson Spearman Kendall
Toufu-vicon 0.799 0.871 0.871
Girls-vicon 0.843 0.865 0.865
In physical imaging formation process, the camera first captures a RAW image $Y(p)$, and the relationship between the RAW image $Y(p)$ and the scene radiance $X(p)$ is as follows:
$X(p)=\frac{Y(p)}{tg\pi\left(\frac{f}{2a}\right)^{2}}$
where $f$ is the focal length, $t$ is the exposure time, $g$ is the gain, $a$ is the aperture f-number. Like most 3D HDR methods[17, 4, 13], the $f$, $g$ and $a$ are fixed but unknown in experiments. Thus, the exposure times that we estimate have a constant scale with the ground truth exposure times. To evaluate the exposure time estimation, we used metrics such as Pearson correlation coefficient, Spearman rank correlation coefficient, and Kendall Tau-b correlation coefficient. A correlation value greater than 0.7 is generally considered a strong correlation. The quantitative results are shown in Table 7, the results show that we can estimate exposure times :
Figure 9 shows the comparison between the estimated exposure times and the ground truth exposure times for each training image in scenes Toufu-vicon (above) and Girls-vicon (bottom). The results are scaled uniformly, and it can be observed that the estimated exposure times closely follow the trend of the ground truth exposure times.
6.3 More experimental results under training view.
Figure 8: Qualitative comparison on synthetic dataset (Trolley, Factory, Cozyroom) under training view. BAD-Gaussians can deblur the training views as ours. However, due to the failure of pose optimization in the BAD-Gaussians, its image are misaligned with others.
The results in Figure 8 demonstrate that our method effectively deblurs images under training views and achieves better image quality compared to other methods. It is worth noting that while BAD-Gaussians [64] is also capable of deblurring images under training view, its lack of robustness to varying brightness conditions leads to pose optimization failure. As a result, The performance of deblurring is poor, even causing misalignment in the images under the training view.
Figure 9: Comparison between estimated exposure times and ground truth exposure times. Toufu-vicon (above) and Girls-vicon (bottom).
Table 8: Quantitative comparisons on the synthetic datasets in term of deblur.
Factory Pool Trolley Cozyroom
[…]
BAD-GS [64] 24.32 0.73 0.12 25.87 0.79 0.23 19.06 0.62 0.19 23.37 0.79 0.11
BAD-GS+bilagrid [56] 28.25 0.79 0.08 31.99 0.86 0.06 22.16 0.65 0.15 26.48 0.81 0.09
CasualHDRSplat-random (ours) 31.20 0.88 0.05 32.95 0.87 0.10 23.65 0.69 0.12 29.60 0.84 0.05
CasualHDRSplat-gt (ours) 32.00 0.91 0.07 34.53 0.96 0.05 29.35 0.87 0.08 33.01 0.93 0.04
Table 9: Quantitative comparisons on the ScanNet dataset in term of deblur (using BRISQUE metric).
scene0024_01 scene0031_00 scene0036_00 scene0072_01 scene0077_00 scene0489_02 Average
BAD-GS [64] 42.78 60.48 57.37 42.72 67.30 65.01 55.94
CasualHDRSplat-random (ours) 38.08 47.65 53.37 37.55 62.35 58.23 49.53
We added comparison against the bilateral grid method [56] applied to gsplat [59] and BAD-Gaussians [64] in Table 8 and Table 13. The bilateral grids applied to NeRFs and 3DGS gives robustness to large appearance changes, enabling high quality 3D LDR reconstruction and mid-tone rendering quality. However, bilateral grids are not compatible with representing a 3D HDR scene, thus gives degraded renderings on high-contrast, over-exposured and under-exposured views.
In addition, we also evaluated the quantitative metrics for deblurring on public real-world datasets, e.g. ScanNet datastes. Due to most images of ScanNet dataset are motion-blurred, we can not find sharp reference images for evaluating, thus we utilize the no-reference image quality metric BRISQUE [38] to quantitatively compare the deblurring performance between our method and BAD-Gaussians [64], as shown in Table 9.
Table 10: Quantitative comparisons on Realsense and SmartPhone dataset under novel view.
Yakitori Toufu
[…]
gsplat [59] 25.04 0.83 0.27 29.88 0.81 0.24
BAD-Gaussians [64] 23.31 0.78 0.28 30.05 0.82 0.24
HDR-Plenoxels [17] 27.13 0.81 0.33 30.91 0.82 0.29
Gaussian-W [62] 27.57 0.84 0.28 30.89 0.83 0.26
CasualHDRSplat-random (ours) 28.56 0.84 0.22 32.75 0.87 0.17
CasualHDRSplat-gt (ours) 29.19 0.87 0.16 32.84 0.91 0.18
Fish-pixel8pro Building-pixel8pro Toufu-vicon Girls-vicon
[…]
gsplat [59] 23.20 0.82 0.16 25.99 0.81 0.11 24.34 0.81 0.28 23.81 0.77 0.28
gsplat+bilagrid [56] 25.26 0.78 0.14 25.47 0.77 0.16 30.48 0.82 0.17 26.76 0.69 0.25
BAD-GS [64] 24.28 0.78 0.14 26.93 0.82 0.11 24.22 0.82 0.24 23.95 0.77 0.28
BAD-GS+bilagrid [56] 25.12 0.77 0.17 25.63 0.77 0.15 30.52 0.83 0.17 26.18 0.71 0.23
HDR-Plenoxels [17] 19.39 0.53 0.65 26.87 0.81 0.15 17.90 0.51 0.69 26.73 0.84 0.30
Gaussian-W [62] 26.13 0.83 0.15 27.99 0.82 0.11 26.38 0.83 0.29 26.88 0.86 0.25
CasualHDRSplat-random (ours) 28.30 0.83 0.13 28.79 0.83 0.09 30.87 0.90 0.15 32.00 0.90 0.19
CasualHDRSplat-gt (ours) 30.81 0.87 0.12 29.71 0.85 0.08 31.34 0.92 0.12 32.39 0.91 0.17
Figure 10: Qualitative comparison for pose estimation on the Girls-vicon sequence of the Realsense dataset.
Figure 11: Qualitative comparison for pose estimation on the Toufu-vicon sequence of the Realsense dataset.
6.4 More experimental results under novel view.
The quantitative experimental results in Table 10 indicate that our method significantly outperforms previous approaches in novel view synthesis on two real-world datasets. Further, the qualitative experimental results in Figure 13 and Figure 14 demonstrate that our method produces higher-quality rendered images under novel viewpoints compared to other approaches. These results indicate that our method is capable of learning accurate HDR scene representations and implicit CRF representations.
Figure 12: Qualitative comparison with bilateral method on Smartphone dataset under novel view. It is better to view the results on a monitor with high resolution and a gamut coverage close or better than sRGB.
In addition, we compare against the bilateral grid method [56] applied to gsplat [59] and BAD-Gaussians [64], as shown in Table 10 and Figure 12. As aforementioned, with bilateral grid [56], BAD-Gaussians [64] cannot represent the HDR details of the 3D scenes, thus yields degraded renderings in the high-contrast areas. As it is showed in the Figure 12, the girls in the Fish sequence of the Smartphone dataset are over-exposed in some views, thus exhibits under-saturation; Meanwhile, the duck in the Building sequence of the Smartphone dataset has lost its details and exhibits artifacts on its edge.
6.5 Video results of CasualHDRSplat
To show the great performance of our approach, we provide two supplementary videos.
novel_view_spiral_traj.mp4 uses the spiral trajectory to render novel views with the given exposure times. The upper half of the video shows the rendered images while the lower half of the video shows the exposure time of current frame, as the camera moves along a spiral trajectory.
train_view_traj.mp4 uses the trajectory and exposure times of training view. On the lower half of the video, we show that our approach is robust to blurry inputs and can render the sharp images with given exposure times.
Figure 13: Qualitative comparison on synthetic dataset(Cozyroom, Factory, Outdoorpool) under novel view.
BAD-Gaussians HDR-Plenoxels Gaussian-W Ours-random Ours-gt Reference
Figure 14: Qualitative comparison on Smartphone dataset under novel view.
6.6 More experimental results on pose estimation.
To demonstrate that our method can accurately recover the continuous camera motion trajectory, we visualized and compared the trajectory optimized by our method with the ground truth trajectory, as well as with other baselines. The qualitative results in Figure 10 and Figure 11 indicate that our method achieves higher pose estimation accuracy compared to previous methods.